# Towards Efficient Training of Graph Neural Networks: A Multiscale Approach

### v2

@@ Appendix B Additional Results for Transductive Learning
Random and subgraphs 76.44% (0.1%) 77.12% (0.3%) 76.92% (0.8%) 79.29% (0.1%) 78.59% (0.5%) 78.28% (0.6%)
GIN Fine grid          1 level: 80.58% (4.0%)          1 level: 71.73% (0.4%)
@@
Random and subgraphs 80.58% (14.7%) 82.12% (3.8%) 81.35% (3.0%) 72.64% (0.4%) 73.71% (0.8%) 74.16% (1.3%)
GAT Fine grid          1 level: 71.44% (3.9%)          1 level: 79.09% (1.1%)
@@
Random and subgraphs 72.98% (3.9%) 74.13% (1.9%) 76.15% (3.1%) 78.93% (0.6%) 79.20% (0.2%) 79.60% (0.9%)
Table 10: Comparison of different methods and coarsening levels for BlogCatalog and WikiCS datasets.
@@
Table 11: Experimental setup for transductive learning datasets.
Dataset Nodes Edges Features Classes
Cora Sen et al. (2008) 2,708 10,556 1,433 7
Citeseer Sen et al. (2008) 3,327 9,104 3,703 6
Pubmed Sen et al. (2008) 19,717 88,648 500 3
Flickr Zeng et al. (2019) 89,250 899,756 500 7
WikiCS Mernyei & Cangea (2020) 11,701 216,123 300 10
DBLP Bojchevski & Günnemanng (2017) 17,716 105,734 1,639 4
Facebook Yang et al. (2020) 4,039 88,234 1,283 193
BlogCatalog Yang et al. (2020) 5,196 343,486 8,189 6
PPI (transductive) Yang et al. (2020) 56,944 1,612,348 50 121
Table 12: Datasets statistics.
Appendix C Additional Results for Shapenet Dataset
@@
To emphasize the efficiency of our methods, we further provide timing analysis of Shapanet ’Airplanes’ data computations. Timing was calculated using GPU computations, and we averaged across 5 epochs after the training loss was stabilized. Results appear in Table 14 for multiscale training and in Table 15 for multiscale gradient computations, where we show that our training process was indeed faster than the original process which only took into consideration the fine grid.
Fine Grid Random Subgraphs Random and Subgraphs
@@
Fine Grid Random Subgraphs Random and Subgraphs
@@
Fine Grid Random Subgraphs Random and Subgraphs
@@
Table 16: Experimental setup for ShapeNet dataset.
Appendix D Additional Details and Results for PPI Dataset

### v3

@@ Appendix E Additional Transductive Learning datasets
  Rand. & Sub. 76.44% (0.1%) 77.12% (0.3%) 76.92% (0.8%) 79.29% (0.1%) 78.59% (0.5%) 78.28% (0.6%)
GIN
Fine Grid ($L=1$) 80.58% (4.0%) 71.73% (0.4%)
@@
  Rand. & Sub. 80.58% (14.7%) 82.12% (3.8%) 81.35% (3.0%) 72.64% (0.4%) 73.71% (0.8%) 74.16% (1.3%)
GAT
Fine Grid ($L=1$) 71.44% (3.9%) 79.09% (1.1%)
@@
  Rand. & Sub. 72.98% (3.9%) 74.13% (1.9%) 76.15% (3.1%) 78.93% (0.6%) 79.20% (0.2%) 79.60% (0.9%)
Table 16: Experimental setup for transductive learning datasets.
@@
Appendix F Additional Results for Shapenet Dataset
@@
To emphasize the efficiency of our methods, we further provide timing analysis of Shapanet ’Airplanes’ data computations. Timing was calculated using GPU computations, and we averaged across 5 epochs after the training loss was stabilized. Results appear in Table 18 for multiscale training and in Table 19 for multiscale gradient computations, where we show that our training process was indeed faster than the original process, which only took into consideration the fine grid. We provide additional experimental details in Table 20.
Graph Density (kNN) Fine Grid Random Subgraphs Rand. & Sub.
@@
Graph Density (kNN) Fine Grid Random Subgraphs Rand. & Sub.
@@
Graph Density (kNN) Fine Grid Random Subgraphs Rand. & Sub.
@@
Appendix G Details and Results for PPI Dataset